# A Novel Spike Transformer Network for Depth Estimation from Event Cameras via Cross-modality Knowledge Distillation

### v2

@@ 4.4 Experiment Result
4.4.1 Qualitative and Quantitative Comparisons
Our experimental investigation encompasses both quantitative performance and energy consumption analyses, utilizing synthetic (DENSE) and real-world (DESC) datasets. A total of nine metrics were employed to evaluate the outcomes comprehensively.
@@ 4.4.1 Qualitative and Quantitative Comparisons
Table 1: Quantitative performance comparison on the synthetic datasets (DENSE) using various models. Results for both the validation set and the test set are presented. Symbols $\downarrow$ and $\uparrow$ indicate that a lower value and higher value are preferable, respectively. The "Param" column refers to the number of parameters, and "Power" provides the average theoretical energy consumption for predicting an image.
Table 1 presents a quantitative performance comparison using synthetic datasets (specifically the DENSE dataset). The results consistently indicate that the proposed method outperforms the alternative approaches across nearly all evaluated metrics. Notably, significant enhancements are observed in the metrics of absolute relative error (Abs.Rel) and squared relative error (Sq.Rel). These are particularly critical metrics in depth estimation tasks. The Abs.Rel for the proposed method recorded at 0.80, significantly surpasses those of U-Net (2.89), E2Depth (9.91), and Spike-T (1.57), representing improvements of 72%, 91.9%, and 49% respectively. Similarly, the Sq.Rel of our method stands at 8.32, substantially lower than those of U-Net (72.25), E2Depth (96.01), and Spike-T (39.77), indicating improvements of 88.5%, 91.3%, and 79% respectively. Additionally, our method demonstrates marginal enhancements in accuracy metrics ($\delta<1.25$, $\delta<1.25^{2}$, and $\delta<1.25^{3}$), achieving scores of 0.53, 0.68, and 0.76, respectively, which are slightly better than the competing methods.
In terms of power consumption, the proposed method, which leverages pure spike computing, shows a marked advantage over its competitors. Furthermore, the adoption of knowledge distillation has facilitated the use of merely four Transformer blocks in our method, significantly reducing the parameters compared to the Spike-T method, which utilizes eight Transformer blocks.
@@ 4.4.2 Ablation study
Figure 10: Visualization results on validation synthetic dataset by using fcn head and proposed fusion depth estimation head.
Figure 11: Visualization results on validation synthetic dataset with and without knowledge distillation.
@@ 5 Discussion
Transformer-based SNNs represent a novel form of SNNs that combines the transformer architecture with SNNs, providing great potential to break through the performance bottleneck of SNNs. Yao et al. [43], and Zhou et al[27], proposed two different Spike-Driven Self-Attention models. They utilised only mask and addition operations to avoid multiplication, which are efficient and have low computational energy consumption. Zhou et al [28], proposed Spikingformer , which modified the residual connection to purely event-driven, which is energy efficient while improving performance. However, the transformer based SNNs is still in its infancy in depth estimation[26]. There are two major issues: 1) the difficulty in training pure transformer based SNN models; 2) and the limited availability of paired depth data in event data to support transformer model training. Knowledge distillation provides a manner to train new or small models using well pretrained models. In this work, we propose a knowledge distillation method to bring ANN model knowledge into SNN.
This work introduces a novel spike transformer network for the monocular depth estimation task from event camera data. Our approach leverages the powerful spatial-temporal modeling capabilities of transformer architectures while adhering to the principles of spiking neural networks (SNNs) through the use of spike-driven residual learning and spike self-attention mechanisms.
Firstly, the proposed SNN backbone of spike transformers, driven purely by spikes, has demonstrated excellent capability and efficiency in extracting features from spike event data, achieving significant improvements in several evaluation metrics, resulting in satisfactory deep prediction results. Meanwhile, by eliminating the need for floating-point and integer-float multiplications, our approach significant energy savings while maintaining strong performance. This energy efficiency is a critical advantage for practical applications of SNNs[53, 54].
Moreover, the ablation study underscores the significance of the proposed fusion depth estimation head and knowledge distillation components. The fusion head, designed to fuse features from multiple transformer stages, proved helpful in preserving spatial and temporal information and recovering fine details in the depth predictions. Meanwhile, our proposed single-stage cross-modality knowledge distillation framework represents a novel and effective strategy for enhancing the performance of SNNs, even when faced with limited training data. By leveraging the knowledge encoded within DINOv2, a large-scale vision foundation model pre-trained on vast amounts of data, our SNN model can effectively capture relevant spatio-temporal features and mitigate the effects of noise and uncertainty inherent in spiking data, thereby improving its generalization capabilities. This approach effectively solves the overfitting problem faced by SNNs trained on small datasets, which seriously hinders the wide application of SNNs in various fields[5, 14].
The experimental results presented in this study provide compelling evidence of the efficacy of our proposed method. On both synthetic (DENSE) and real-world (DESC) event camera datasets, our model consistently outperformed existing state-of-the-art approaches, demonstrating substantial improvements in critical metrics such as absolute relative error and squared relative error. The visualization of depth predictions further highlights the model’s ability to capture intricate details, sharp edges, and contours, even in challenging low-light conditions.
While the results presented in this work are promising, further experiments and improvements are needed. For instance, deploying and evaluating the proposed SNN on neuromorphic computing hardware platform such as SpiNNaker[55], BrainScales[56] and TrueNorth[57]. As hardware acceleration for SNNs continues to advance, it will be crucial to assess the real-world performance and energy efficiency of our model in practical scenarios, such as autonomous vehicle navigation or robotic perception tasks.
It is also worth noting that while our work focused specifically on the task of depth estimation, the proposed spike transformer architecture and knowledge distillation framework could potentially be extended to other computer vision tasks. The inherent property of event camera data to capture high-rate temporal features provides an opportunity to utilise our proposed SNN in other dynamic scene tasks such as object tracking, motion estimation or event prediction. By leveraging the cross-modality knowledge distillation from the large foundation models and the powerful feature capture capabilities of transformer architecture, the proposed spike transformer has the ability to extract valuable features for the temporal evolution of complex environments, enabling more intelligent and reactive systems. This remains an attractive direction for future research.
@@ 6 Conclusion
This paper presents a novel Spike Transformer network designed for depth estimation using data from spiking cameras. By integrating spike-driven residual learning and spiking self-attention mechanisms, we have developed a transformer architecture driven entirely by spikes. This design markedly increases the computational efficiency of Spiking Neural Networks (SNNs). Moreover, our single-stage knowledge transfer framework, which draws on large foundational ANN models such as DINO V2, enhances the performance of SNNs even with a limited data availability. Our experimental evaluations on both synthetic and real datasets demonstrate significant improvements across several metrics, notably in Abs Rel and Sq Rel (49% and 39.77% improvements over the state-of-the-art model Spike-T, respectively). In addition to its accuracy, the proposed model also exhibits lower power consumption, a critical attribute for practical applications. Future research will focus on further validation and deployment on real datasets and dedicated SNN processors, thus potentially broadening the applicability of Spike Transformers in real-world scenarios.
Acknowledgements (not compulsory)
This work is supported by the BBSRC (Projects BB/R019983/1, BB/S020969/1 and EP/X013707/1).
Author contributions statement
X.Z. and LX.H. contributed to the conception and design of the study, while X.Z. and T.S. conducted the experiment(s). X.Z., LX.H., T.S., LH.H., and D.D. contributed to the analysis. LX.H. managed the project, acquired funding, and provided supervision. All authors contributed to the drafting and revision of the manuscript.

### v3

@@ 4.4 Experiment Result
4.4.1 Model performances
Our experimental investigation encompasses both quantitative performance and energy consumption analyses, utilizing synthetic (DENSE) and real-world (DESC) datasets. Several metrics were employed to evaluate the outcomes comprehensively.
@@ 4.4.1 Model performances
Table 1: Quantitative performance comparison on the synthetic datasets (DENSE) using various models. Results for both the validation set and the test set are presented. Symbols $\downarrow$ and $\uparrow$ indicate that a lower value and higher value are preferable, respectively. The ”Param” column refers to the number of parameters, and ”Power” provides the average theoretical energy consumption for predicting an image.
Table 1 presents a quantitative performance comparison using synthetic datasets (specifically the DENSE dataset).The results consistently indicate that the proposed method outperforms the alternative approaches across nearly all evaluated metrics.
Notably, substantial reductions are observed in the absolute relative error (Abs.Rel) and squared relative error (Sq.Rel)—critical metrics in depth estimation tasks. The Abs.Rel for the proposed method, recorded at 0.80, represents reductions of 72%, 91.9%, and 49% compared to U-Net (2.89), E2Depth (9.91), and Spike-T (1.57), respectively. Similarly, the Sq.Rel of our method, at 8.32, demonstrates reductions of 88.5%, 91.3%, and 79% relative to U-Net (72.25), E2Depth (96.01), and Spike-T (39.77).
In addition to error reduction, the proposed method achieves modest increases in accuracy metrics ($\delta<1.25$, $\delta<1.25^{2}$, and $\delta<1.25^{3}$), attaining values of 0.53, 0.68, and 0.76, respectively. These values are slightly higher than those achieved by the competing methods, reinforcing the method’s superior capability in depth estimation tasks.
In terms of power consumption, the proposed method, which leverages pure spike computing, shows a marked advantage over its competitors, reducing power consumption by up to 82.9% compared to U-Net (from 72.93mJ to 12.43mJ). Furthermore, the adoption of knowledge distillation has facilitated the use of merely four Transformer blocks in our method, significantly reducing the parameters by 42.4% compared to the Spike-T method (from 35.68M to 20.55M), which utilizes eight Transformer blocks.
@@ 4.4.2 Ablation study
Figure 10: Visualization results on validation synthetic (DENSE) dataset by using FCN head and proposed fusion depth estimation head.
Figure 11: Visualization results on validation synthetic (DENSE) dataset with and without knowledge distillation.
@@ 5 Discussion
The development of efficient and accurate depth estimation methods for event cameras remains a significant challenge in computer vision(Gallego et al., 2022; Tayarani-Najaran and Schmuker, 2021). Our work addresses this challenge through three key innovations: a purely spike-driven transformer architecture, fusion depth estimation head and a novel knowledge distillation framework. The experimental results demonstrate the improvements in both accuracy and energy efficiency, we discuses the implications, limitations, and potential impact of our approach in this section.
The satisfactory performance of the spike transformer architecture can be attributed to several factors. First, by implementing spike-driven residual learning and self-attention mechanisms (Zhou et al., 2022, 2023), our model effectively captures the temporal dynamics inherent in event camera data while maintaining the computational efficiency characteristic of SNNs (Rafi, 2021). This is evidenced by the significant reductions in absolute relative error (49% reduction) and squared relative error (39.77% reduction) compared to the state-of-the-art Spike-T model (Zhang et al., 2022). The elimination of floating-point operations in the transformer portion substantially reduces power consumption (Han and Roy, 2020; Lemaire et al., 2022), making our approach more practical for real-world applications where energy efficiency is crucial.
Our innovative fusion depth estimation head integrates multi-scale transformer features, preserving fine details and global structure, addressing limitations of traditional methods and ensuring robust performance. Traditional approaches often struggle to maintain fine-grained spatial information while processing temporal event data (Ming et al., 2021). Our fusion head addresses this by effectively combining features from multiple transformer stages, as demonstrated by the improved preservation of detail in our depth predictions compared to standard linear FCN heads (Ronneberger et al., 2015). This multi-scale feature integration is particularly beneficial for capturing both fine details and global scene structure (Fan et al., 2021), contributing to the overall robustness of our depth estimates. However, it is important to acknowledge the limitation that the fusion depth estimation head is not purely spike-based. This decision was driven by the requirement for high accuracy in depth estimation tasks. Pure spike-based operations, while more energy-efficient, currently face limitations for precision value prediction.
Our hybrid approach allows us to leverage the energy efficiency of spike-based computing in the feature extraction stages while maintaining the high accuracy requirements of depth estimation through conventional computation in the fusion head.
Our knowledge distillation framework represents a significant advancement in addressing the longstanding challenge of training SNNs with limited data (He et al., 2023). By leveraging the rich feature representations learned by DINOv2 (Oquab et al., 2023) from its vast training dataset, our approach effectively transfers knowledge across modalities (from traditional images to event data). Ablation studies clearly demonstrate the value of this knowledge transfer, showing marked improvements in depth estimation accuracy and robustness to noise when knowledge distillation is employed (Gou et al., 2021). This is particularly evident in the model’s ability to handle challenging scenarios such as estimating depths of distant objects and operating in low-light conditions (Furmonas et al., 2022).
Several avenues for future research emerge from our findings. First, investigating methods to achieve comparable accuracy with pure spike-based fusion mechanisms remains an important challenge, though this may require fundamental advances in spike-based computing precision. Second, evaluation on neuromorphic hardware platforms (such as SpiNNaker 1/2 (Furber et al., 2014; Mayr et al., 2019), BrainScales (Pehle et al., 2022), or TrueNorth (Löhr et al., 2020)) would provide valuable insights into real-world performance and energy efficiency. Especially SpiNNaker 2(Mayr et al., 2019), unlike traditional methods, it is specifically designed to handle operations such as multiplications, which are generally inefficient for spiking computations. The introduction of such platforms highlights the necessity for hybrid approaches, where specific operations may leverage conventional hardware optimizations while retaining the efficiency of spike-driven designs.
Additionally, our architecture’s ability to effectively process temporal event data suggests potential applications beyond depth estimation, such as object tracking (Chen and Guo, 2019) or motion estimation (Huang et al., 2022). The success of our knowledge distillation approach also raises interesting questions about the broader applicability of foundation models in training efficient SNNs for various computer vision tasks (Bommasani et al., 2021).
The integration of event cameras in autonomous systems and robotics applications continues to grow (Cao et al., 2021), driven by their advantages in terms of latency, dynamic range, and power efficiency (Posch et al., 2011). Our work demonstrates that by combining the biological inspiration of SNNs with modern deep learning architectures and knowledge distillation techniques, we can develop more efficient and accurate methods for processing event camera data. While our hybrid approach represents a careful balance between computational efficiency and accuracy requirements, it provides valuable insights for the broader development of neuromorphic computing systems (Auge et al., 2021). The success of this approach suggests that future developments in spike-based computing may benefit from similar pragmatic trade-offs between pure neuromorphic computation and task-specific performance requirements.
@@ 6 Conclusion
In this paper, we have introduced a novel energy-efficient Spike Transformer network for depth estimation, leveraging spiking camera data. The proposed architecture integrates spike-driven residual learning and spiking self-attention mechanisms, creating a transformer framework that operates entirely within the spike domain. This innovative design achieves significant computational efficiency, with an 82.9% reduction in power consumption compared to conventional methods (from 72.93mJ to 12.43mJ per inference). Additionally, our single-stage knowledge distillation framework, leveraging large foundational ANN models such as DINOv2, enables robust training of Spiking Neural Networks (SNNs) even in the presence of limited data. Extensive evaluations on synthetic and real-world datasets demonstrate the efficacy of our approach, with significant improvements in key performance metrics, including a 49% reduction in Absolute Relative Error and a 39.77% reduction in Square Relative Error compared to the state-of-the-art SpikeT model. The architecture further enhances efficiency with a 42.4% reduction in parameters (20.55M versus 35.68M), making it particularly well-suited for resource-constrained environments. By combining high accuracy with remarkably low power requirements, our spike-based design is well-suited for practical applications. Future work will focus on extending this research to broader real-world scenarios, including deployment on dedicated SNN processors and further validation with diverse datasets. These efforts aim to unlock the full potential of Spike Transformers in applications such as autonomous navigation, robotics, and energy-efficient vision systems, paving the way for advanced neuromorphic computing in practical settings .
Acknowledgement
Funding acknowledgment: EPSRC (EP/X013707/1), BBSRC(BB/R019983/1, BB/S020969/1), Innovate UK(10091423).